On January 1, 2019, Crane Company purchased furniture for $7830. The company expects to use the furniture for 3 years. The asset has no salvage value. The book value of the furniture at December 31, 2020 is

a

$2610.

b

$5220.

c

$7830.

d

$0.
Answer #1

a) $2610

Initial Cost : $7830

Estimated Useful Life: 3 years

Salvage Value : $0

Depreciation rate = 1/3 *100 = 33%

Annual Depreciation = 7830*33% = $2,610 = $2,610 *2 = $5,220

Book Value as on 31 December 2020 = $7,830 - $5,220 = $2,610

Dear Student, if you like the answer be sure to give a thumbs up!